# Deep-learning-based continuous attacks on QKD protocols

### v2

@@ III.3 Attack via continuous measurement
We now discuss our new kind of attacks, based on an homodyne measurement of the photon that is fed to a LSTM neural network. When modelling the dynamics of photons under homodyne detection, one must set the measurement operator $e$ of Eq. (3). First, let us use
@@
and consider in the first instance that the homodyne measurement is performed during the whole travel time, set to $\gamma_{D}t_{f}=3$, and with other parameters $\eta=0.5$, $\omega=\gamma_{E}=\gamma_{D}$. With these parameters we get, from the solutions of Eq. (3), a QBER of $49\%$ (red line in Fig 3), much higher than the $37\%$ of the projective measurement obtained from Eq. (10) and the $24.4\%$ of the case with dissipation only, obtained from Eq. (8). In addition, we get a neural network test accuracy $A\approx 71\%$, which is within the interval given by Eq. (9). We interpret this result as coming from the fact that the states |$\ket{0}$ and |$\ket{1}$ being eigenstates of the measurement operator $e=\sigma_{z}$, the neural network can approximately distinguish these two while it must do random guess on the other two states ($\ket{+}$ and $\ket{-}$), which would lead to a test accuracy around $75\%$ (see Appendix D). This suggests that the neural network puts an heavy weight on the early-time behavior of the photocurrent, as confirmed in Appendix E. Hence, we see that the spying accuracy achieved via this simple continuous measurement scheme is comparable with the one achieved via the projective measurement, but the QBER is higher.
@@
Let us first look at the neural network test accuracy $A$ of this new measurement operator as a function of $\theta/\pi$, which is depicted in Fig. 4, together with the associated standard deviation. There are four angles leading to an accuracy around $90\%$, as summarized in Table 1, which is much higher than the $71\%$ found earlier. Note that the four angles seem equivalent given their values and the standard deviations.
@@
To obtain the impact of the measurement on the BB84 protocol itself we average the QBER over the four possible initial states, and evaluate it as a function of $\theta$ and $\gamma_{D}t$, as displayed in Fig. 5.
@@
One can see there is a trade-off between the amount of information we can extract and the perturbation that this measurement induces. Indeed, the angles that minimize the QBER, which are $\theta=0\pm k\pi,k\in\mathbb{Z}$, yield a low test accuracy, thus decreasing the information the spy obtains (see Fig. 4). In order to quantify this tradeoff, we define a new quantity $\lambda(\theta)$ as the QBER divided by the accuracy $A$ of the network for a given measurement basis (i.e., a given $\theta$)
@@
which is shown for $\gamma_{D}t_{f}=3$ in Fig. 6. We see that among the four measurement angles maximizing the spy accuracy (circled in red), $\theta=1.86\pi$ yields the lowest $\lambda$ ratio, with an accuracy arround $90\%$ and a QBER around $37\%$, though the other angles give similar performances.
@@
The top panel of Fig. 3 gathers all the QBER discussed above. In addition, three other lines are present: the QBER for a continuous measurement when $\theta$ is set to minimize the QBER itself (green line), to maximize the accuracy (orange line), and finally to minimize $\lambda$ (blue line). In the bottom panel, the $\theta$ values of these three lines are represented, along with the four optimal angles from Table 1 as guides to the eye (black dashed lines). The QBER optimized with respect to itself is at first higher than the one from no measurement at all, before converging to the same value of $25\%$. However, even though Alice and Bob could never detect this eavesdropping (at least for large $\gamma_{D}t_{f}$), the neural network manages to learn something from it since its mean accuracy for this case ($\theta=0$, i.e., $e=\sigma_{x}$) is higher than $25\%$ (see Fig. 4). Also, the measurement optimized with respect to the accuracy yields a lower QBER than a projective measurement for $\gamma_{D}t_{f}\leq 2.5$ and the one optimized with respect to $\lambda$ yields an even lower QBER, but still higher than the green line.
Finally, to make the measurement more realistic, one can reduce the duration of the continuous measurement so that it starts later than the initial time and finishes before the final time. We choose here the optimal measurement angle found earlier, $\theta=1.86\pi$, and set it to start at $\gamma_{D}t=0.1$ and define a duration $\Delta t$ of the continuous measurement with the aim of decreasing its impact on the qubit states while keeping a reasonable neural network test accuracy. As shown in Fig. 7, which displays $A$ as a function of $\gamma_{D}\Delta t$, the accuracy reaches $86.1\%$ with a reasonable standard deviation ($2\%$) when measuring between $\gamma_{D}t=0.1$ and $0.5$. For longer duration, we see that the accuracy is reasonable but the standard deviation becomes very high as compared to Fig. 4. This suggests that the initial time-evolution (i.e., before $\gamma_{D}t=0.1$), which is not fed to the neural network here, is crucial to make it sharp. Altogether, taking $\theta=1.86\pi$ and the optimal duration $\gamma_{D}\Delta t=0.4$, we obtain an accuracy of $86.1\%$ and a QBER of $27.6\%$, this latter representing only a $2.6\%$ increase compared to the time-evolved state where no measurement is made. Comparatively, the Time shift [14, 15] attack scheme yields an accuracy around $60-70\%$ and an increase of the QBER between $1$ and $2\%$. Hence, though the noise model for the fiber is not the same, our new attack scheme seems to provide similar QBER increases and very high accuracies, which suggests it is a potential threat that should be investigated further. A summary of the results of this section can be found in Table 2.
@@ III.4 Attack via continuous measurement and quantum feedback
As a final section, we briefly investigate here the possibility for the spy to apply quantum feedback based on their measurement [27] with the aim of covering their track. We thus adapt Eq. (3) to include the action of a feedback which yields [27]
@@
$+\gamma_{E}\mathcal{D}[f]\rho_{J}\frac{dt}{\eta}+\gamma_{E}dW\mathcal{H}\left[\sqrt{\eta}e-if/\sqrt{\eta}\right]\rho_{J}$
@@
Dissipation Attack Impact (QBER) Information gain (Accuracy)
@@
IV Conclusion
@@
In this paper, we introduced a new type of attack on QKD protocols based on continuous measurement that, used as an input of a trained recurrent neural network, allows the spy to retrieve with high accuracy the initial states sent by one of the parties without being significantly noticed. We denotes our attack as a Deep-learning-based continuous attack (DLCA). Although more quantitative and comparison analyses should be done, also in terms of noise models considered for the optical fiber, our attack scheme exhibits performances similar to other attacks, which makes it a potential serious threat worthy to be explored in future works.
@@
To go further, one could use a more complex and realistic noise model of the optical fiber. In [24], Kozubov et al. span the space using three states: the vacuum state and the states we denoted |$\ket{0}$ and |$\ket{1}$ in this work (i.e., horizontally and vertically polarized photons). By doing so, they take into account the non-zero probability that the photon is absorbed in the optical fiber. They also tune the phenomenological parameters involved in the master equation, which allows to take into account the phenomena of birefringence, isotropic absorption, and dichroism. However, the depolarizing quantum channel [30] is a commonly used noise model and would be worth considering in our study. In addition, one could consider the potential losses caused by the imperfection of Alice and Bob detectors.
@@
In addition, one could investigate practical implementations of our attack scheme involving homodyne detection of single photon [48], by exploiting evanescent waves in optical fibers [49] or quantum memories [50].
@@
The next step could also be to determine the possible countermeasures one could implement to make the BB84 protocol robust against our attack, such as decoy states protocols [9, 51]. However, one could also investigate how our scheme could be applied to coherent states continuous variable protocols [52, 53] exploiting homodyne or heterodyne detection, or entanglement-based protocols such as the E91 [54], the BBM92 [55] or on device-independent protocols [56, 57]. Entanglement-based protocols are promising for satellites QKD, which is currently being extensively studied by the scientific community [58, 59, 60, 61]. One could thus explore the generalization of our attack and its practical implementation to satellite QKD.
@@ Acknowledgements.
We thank Jérôme Denis and John Martin for helpful initial discussions on the topic.
Appendix A Analytical derivation of the QBER without attacks

### v3

@@ III.3 Attack via continuous measurement
We now discuss our new kind of attack, based on an homodyne measurement of the photon that is fed to a LSTM neural network. When modeling the dynamics of photons under homodyne detection, one must set the measurement operator $e$ of Eq. (3). First, let us use
@@
and consider in the first instance that the homodyne measurement is performed during the whole travel time, set to $\gamma_{D}t_{f}=3$, and with other parameters $\eta=0.5$, $\omega=\gamma_{E}=\gamma_{D}$. With these parameters we obtain, from the solutions of Eq. (3), a QBER of $49\%$ (red line in Fig 3), much higher than the $37\%$ of the projective measurement obtained from Eq. (10) and the $24.4\%$ of the case with dissipation only, obtained from Eq. (8). In addition, we get a neural network test accuracy $A\approx 71\%$, which is within the interval given by Eq. (9). We interpret this result as coming from the fact that the states |$\ket{0}$ and |$\ket{1}$ being eigenstates of the measurement operator $e=\sigma_{z}$, the neural network can approximately distinguish these two while it must do random guess on the other two states ($\ket{+}$ and $\ket{-}$), which would lead to a test accuracy around $75\%$ (see Appendix E). This suggests that the neural network puts a heavy weight on the early-time behavior of the photo current, as confirmed in Appendix F. Hence, we see that the spying accuracy achieved via this simple continuous measurement scheme is comparable to the one achieved via the projective measurement, but the QBER is higher.
@@
Let us first look at the accuracy $A$ yielded by this new measurement operator as a function of $\theta/\pi$, which is depicted in Fig. 4, together with the associated standard deviation. There are four angles leading to an accuracy around $90\%$, as summarized in Table 1, which is much higher than the $71\%$ found earlier. Note that the four angles seem equivalent given their values and the standard deviations.
@@
To obtain the impact of the measurement on the BB84 protocol itself, we average the QBER over the four possible initial states, and evaluate it as a function of $\theta$ and $\gamma_{D}t$, as displayed in Fig. 5.
@@
One can see there is a trade-off between the accuracy and the disturbance that the measurement induces. Indeed, the angles that minimize the QBER, which are $\theta=0\pm k\pi,k\in\mathbb{Z}$, yield a low test accuracy, thus decreasing the spy accuracy (see Fig. 4). In order to quantify this tradeoff, we define a new quantity $\lambda(\theta)$ as the QBER divided by the accuracy $A$ of the network for a given measurement basis (i.e., a given $\theta$)
@@
which is shown for $\gamma_{D}t_{f}=3$ in Fig. 6. We see that among the four measurement angles maximizing the spy accuracy (circled in red), $\theta=1.86\pi$ yields the lowest $\lambda$ ratio, with an accuracy around $90\%$ and a QBER around $37\%$, though the other angles give similar performances.
@@
The top panel of Fig. 3 gathers all the QBER discussed above. In addition, three other lines are present: the QBER for a continuous measurement when $\theta$ is set to minimize the QBER itself (green line), to maximize the accuracy (orange line), and finally to minimize $\lambda$ (blue line). In the bottom panel, the $\theta$ values of these three lines are represented, along with the four optimal angles from Table 1 as guides to the eye (black dashed lines). The QBER optimized with respect to itself is at first higher than the one from no measurement at all, before converging to the same value of $25\%$. On the lower panel, we see the angle $\theta^{*}_{qber}$ is always equal to 00. This is due to the fact that it corresponds to $e=\sigma_{x}$, which would act in a similar way as the intrinsic dissipation, and thus make the state converge faster to the same steady state yielding a QBER of $25\%$, which is the same value as with no measurement, thus the minimal value a measurement can yield. However, even though Alice and Bob could not in average detect this eavesdropping (at least for large $\gamma_{D}t_{f}$), the neural network manages to learn something from it since its mean accuracy for this case ($\theta=0$, i.e., $e=\sigma_{x}$) is higher than $25\%$ (see Fig. 4). Also, the measurement optimized with respect to the accuracy yields a lower QBER than a projective measurement for $\gamma_{D}t_{f}\leq 2.5$ and the one optimized with respect to $\lambda$ yields an even lower QBER, but still higher than the green line.
Finally, to make the measurement more realistic, one can reduce the duration of the continuous measurement so that it starts later than the initial time and finishes before the final time. We choose here the optimal measurement angle found earlier, $\theta=1.86\pi$, and set it to start at $\gamma_{D}t=0.1$ and define a duration $\Delta t$ of the continuous measurement with the aim of decreasing its impact on the qubit states while maintaining a reasonable accuracy. As shown in Fig. 7, which displays $A$ as a function of $\gamma_{D}\Delta t$, the accuracy reaches $86.1\%$ with a reasonable standard deviation ($2\%$) when measuring between $\gamma_{D}t=0.1$ and $0.5$. For longer duration, we see that the accuracy is reasonable but the standard deviation becomes very high as compared to Fig. 4. This suggests that the initial time-evolution (i.e., before $\gamma_{D}t=0.1$), which is not fed to the neural network here, is crucial to make it sharp. Altogether, taking $\theta=1.86\pi$ and the optimal duration $\gamma_{D}\Delta t=0.4$, we obtain
$A=86.1\%,$ (14)
$\mathrm{QBER}=27.6\%,$ (15)
the latter representing only a $2.6\%$ increase compared to the time-evolved state where no measurement is made. Comparatively, the Time shift [16, 17] attack scheme yields an accuracy around $60-70\%$ and an increase in QBER between $1$ and $2\%$. Hence, though the noise model for the fiber is not the same, our new attack scheme seems to provide similar QBER increases and very high accuracies. A summary of the results of this section can be found in Table 2.
@@ III.4 Attack via continuous measurement and quantum feedback
We briefly investigate here the possibility for the spy to apply quantum feedback based on their measurement [51] with the aim of covering their track. We thus adapt Eq. (3) to include the action of a feedback which yields [51]
@@
$+\gamma_{E}\mathcal{D}[f]\rho_{J}\frac{dt}{\eta}+\gamma_{E}\mathcal{H}\left[\sqrt{\eta}e-if/\sqrt{\eta}\right]\rho_{J}dW$
@@
Dissipation Attack Impact (QBER) Accuracy
@@
IV Information gain and key rates
So far, we have presented the performances of our attack in terms of QBER and accuracy $A$. We now discuss how our attack fits into the existing security proofs and how it compares to optimal individual attacks in terms of information gain. Finally, we close the section by evaluating the typical key rates our attack yields.
IV.1 Information-Disturbance Principle
As predicted by the laws of quantum mechanics, it is impossible to gather information about the identity of a quantum system’s state (when prepared in one of a set of non-orthogonal states) without introducing disturbance in said system [76]. From this, the information-disturbance principle establishes a trade-off between information gained from a measurement and the disturbance caused. For the BB84 protocol, Shor and Preskill suggested in [30] that the maximum (Shannon) information Eve can have about the final key, per bit before privacy amplification, is
$I_{Eve}\leq H_{2}(e_{x})+H_{2}(e_{z}),$ (19)
where $e_{x}$ ($e_{z}$) is the bit (phase) error rate, i.e., the rate of errors in the Z-basis (X-basis), and
$H_{2}(p)=-p\log_{2}(p)-(1-p)\log_{2}(1-p)$ (20)
is the binary entropy function.
An estimator of $I_{Eve}$ is the information gain, or equivalently the expected mutual information, which is defined for two continuous random variables $X$ and $Y$ as
[…]
Let $S$ be a discrete variable representing the different initial states ($s=0,1,2,3$) and $X\in\mathbb{R}^{40}$ a $40$-dimensional continuous variable representing the values of the homodyne currents at each time step, we obtain (see Appendix G) that the information gain yielded by the homodyne measurement on the qubits of the BB84 protocol is
$I(S;x)$ $=0.2894\text{ nats}=0.4175\ \text{bits}.$ (22)
At the same time, the estimated bit and phase error rates being, from Eq. (15), $e_{x}=0.275$ and $e_{z}=0.275$, Eq. (19) becomes
$I_{Eve}\leq 1.7\ \mathrm{bits},$ (23)
showing that Eq. (22) is below the known threshold. Note that this latter is above the $1$ bit of entropy in a single key bit for our specific choice of parameters, which means that the final key rate of the protocol would be negative, such that no secure key can be distilled under standard BB84 security assumptions. This is a consequence of the QBER being above the $11\%$ threshold obtained by Shor and Preskill [30], i.e., the threshold above which the security of the protocol cannot be guaranteed.
If we now consider a regime below the $11\%$ threshold and analyze the information gained by the homodyne measurement when no dissipation is occurring in the optical fiber (such that all of the QBER is caused by the spy), we obtain (with $\gamma_{D}=0$, $\gamma_{E}=\omega/3$ and $2\times 10^{4}$ trajectories)
$\mathrm{QBER}=5.8\%\leq 11\%,$ (24)
$A=88.8\%,$ (25)
$I(S;x)=0.429\ \mathrm{bits}\approx I_{Eve}$ (26)
$I_{Eve}\leq H_{2}(0.058)+H_{2}(0.058)=0.64\ \mathrm{bits}.$ (27)
As $I(S;x)\leq 0.64\ \mathrm{bits}$, this shows that the DLCA still respects the established theoretical bounds.
Comparison with an optimal individual attack.
So far, we have compared our attack to the unconditional bound [Eq. (19)]. Here, we restrict Eve’s power to unitary individual attacks only, for which the information-disturbance principle reads [54]
$I(A;E)\leq\frac{1}{2}\phi\left[2\sqrt{\mathrm{QBER}(1-\mathrm{QBER})}\right],$ (28)
where $I(A;E)$ is the mutual information between Alice and Eve and $\phi[z]\equiv(1+z)\ln(1+z)+(1-z)\ln(1-z)$. It has been shown that such a bound can be saturated using a phase-covariant quantum cloner, which is an approximate cloning procedure for two-level systems on the equator of the Bloch sphere [53], and that a secure key can be distilled as long as the QBER is below $14.65\%$, point at which the curves $I(A;E)$ and $I(A;B)$ intersect. Inserting this value into Eq. (28), we obtain that Eve’s maximum obtainable information is $0.399$ bits. The analysis above showed slightly better performances for the DLCA (even with a lower QBER): 0.4175 bits with dissipation [Eq. (22)] and 0.429 bits without dissipation [Eq. (26)]. This could be explained by the non-unitary nature of our attack.
IV.2 Key rates
The key rate of a QKD protocol is defined as the percentage of secure key bits which can be extracted from the sifted key. It basically makes it possible to calculate the amount of bits Alice and Bob must sacrifice in the error correction and privacy amplification procedure in order to obtain a secure key. Devetak and Winter, in [31], demonstrated the following general and composable bound on the key rate of QKD protocols
$R\geq I(A;B)-I(A;E),$ (29)
where the first term quantifies the error correction cost while the second one quantifies how much privacy amplification is needed. Since the mutual information between two random variables $X$ and $Y$ can be expressed as
$I(X;Y)=H_{2}(X)-H_{2}(X\mid Y)=H_{2}(Y)-H_{2}(Y\mid X),$ (30)
we obtain
$R\geq H_{2}(A\mid E)-H_{2}(A\mid B),$ (31)
which is saturated in the asymptotic limit on infinitely long keys [77]. Below, we evaluate the typical key rates our attack yields under a more realistic noise model: the depolarizing channel model.
Depolarizing channel model.
So far in this paper we have considered a toy model for the intrinsic dissipation occurring in the optical fiber, in order notably to understand the effect of anisotropic dissipation. However, the results presented here could be straightforwardly generalized to more realistic and complex noise models. One such model is the depolarizing quantum channel, an isotropic noise model often used as a simple and effective way to represent noise in quantum communication [56]. For a single qubit, the depolarizing channel is, in a Lindblad form,
$\mathcal{D}_{\mathrm{depol}}(\rho)=\frac{\gamma_{D}}{3}\left(\mathcal{D}\left[\sigma_{x}\right](\rho)+\mathcal{D}\left[\sigma_{y}\right](\rho)+\mathcal{D}\left[\sigma_{z}\right](\rho)\right).$ (32)
DLCA attack.
By using the dissipator (32) with $\gamma_{D}=\omega/100$ and taking the optimal angle and measurement duration found above, Eve’s neural network accuracy about Alice’s key becomes $87.8\%$ and the QBER $10.9\%$. Thus, the conditional entropies, per bit, of Alice’s bit given Eve’s and Bob’s are
$H_{2}(A\mid E)$ $=H_{2}(0.878)=0.535\ \mathrm{bits},$ (33)
$H_{2}(A\mid B)$ $=H_{2}(1-0.109)=0.497\ \mathrm{bits},$ (34)
which yields a final key rate of
$R=0.038\ \mathrm{bits}.$ (35)
Eq. (35) constitutes an upper bound on the usable key rate for Alice and Bob. Indeed, if Alice and Bob make the most pessimistic assumption that the whole QBER of $10.9\%$ is generated by eavesdropping, the key rate they obtain is $0.006$ bits.
Time-shift attack.
For comparison, the Time-Shift attack mentioned above would give
$R=H_{2}(0.7)-H_{2}(0.067)=0.57\ \mathrm{bits},$ (36)
with an accuracy of $70\%$ and a QBER increase of $1\%$ (best values in Table 2). These results show that while our attack does not question the security of the protocol, it does affect the minimum number of resources (and thus the maximum key rate) Alice and Bob could use to secure the protocol.
V Conclusion
@@
In this paper, we introduced a new type of individual attack on QKD protocols based on continuous measurement that, used as an input of a trained recurrent neural network, allows the spy to retrieve with high accuracy the initial states sent by one of the parties without being significantly noticed. We denote our attack as a Deep-learning-based continuous attack (DLCA). Although more quantitative and comparison analyses should be done, also in terms of noise models considered for the optical fiber, our attack scheme exhibits performances similar to the optimal individual attack.
To go further, one could investigate the possible generalization of our strategy to a collective or coherent attack [78].
@@
Also, a more complex and realistic noise model of the optical fiber could be used. In [26], Kozubov et al. span the space using three states: the vacuum state and the states we denoted |$\ket{0}$ and |$\ket{1}$ in this work (i.e., horizontally and vertically polarized photons). By doing so, they take into account the non-zero probability that the photon is absorbed in the optical fiber. They also tune the phenomenological parameters involved in the master equation, which allows to take into account the phenomena of birefringence, isotropic absorption, and dichroism. One could also consider the potential losses caused by the imperfection of Alice and Bob detectors. Overall, it is straightforward to adapt our approach to such other noise models. Note that our attack does not compromise the security of the BB84 protocol as long as Alice and Bob perform enough privacy amplification, reducing the key rate at least below the upper bound we computed.
@@
In addition, one could investigate practical implementations of our attack scheme involving homodyne detection of single photon [79], by exploiting evanescent waves in optical fibers [80] or quantum memories [81], as we partially started in Sec. A.
@@
One could also investigate how our scheme could be applied to decoy states protocols [9, 82], coherent states continuous variable protocols [83, 84] exploiting homodyne or heterodyne detection, or entanglement-based protocols such as the E91 [85], the BBM92 [86] or on device-independent protocols [87, 88]. Entanglement-based protocols are promising for satellites QKD, which is currently being extensively studied by the scientific community [89, 90, 91, 92]. One could thus explore the generalization of our attack and its practical implementation to satellite QKD.
@@ Acknowledgements.
We thank Jérôme Denis and John Martin for helpful initial discussions on the topic. More particularly, we express our sincere thanks to Jérôme Denis for asking the question that inspired this work and for his valuable assistance in the elaboration of the figures. We also thank Daniel Oi for his insightful inputs on information-theoretics-QKD.
Appendix A Homodyne detection schemes
In this section, we derive the stochastic master equation (3) starting from the one modeling the homodyne detection of a damped ancillary field that couples to the photons in the optical fiber, via e.g. their evanescent waves [80].
We consider the stochastic master equation for the full density operator $\rho$ of the combined system made of an optical fiber photon and an ancillary field of the form
$d\rho=$ $-i\left[H+\omega_{a}a^{\dagger}a+ig(ea^{\dagger}-ae^{\dagger}),\rho\right]dt+\gamma_{D}\mathcal{D}[d]\rho dt$ (37)
$+\gamma_{a}\mathcal{D}[a]\rho dt+\sqrt{\gamma_{a}\eta}\mathcal{H}[a]\rho dW,$
where $a$ ($a^{\dagger}$) is the annihilation (creation) operator for the ancillary field of frequency $\omega_{a}$ damped with a rate $\gamma_{a}$. The Heisenberg equation of motion for $a$ reads
$\dot{a}=-(i\omega_{a}+\gamma_{a})a+ge.$ (38)
For $\gamma_{a}\gg g,\omega_{a},\gamma_{D}$, the ancillary field remains weakly populated and can be adiabatically eliminated, as in the bad cavity limit in cavity/circuit QED. According to this, the state of the ancillary field relaxes rapidly and we can set the left-hand-side of the equation above to zero. This makes it possible to slave the ancillary field to the photonic degrees of freedom:
$a\approx\frac{g}{\gamma_{a}}e.$ (39)
Replacing $a$ in Eq. (37) by Eq. (39) directly yields the stochastic master equation (3) of the main text with $\gamma_{E}=g^{2}/\gamma_{a}$.
An alternative implementation of the attack would consist in tapping a small fraction of the quantum signal using a low-reflectivity beam splitter, or, in a photonic integrated circuit (PIC) scenario, a directional coupler. The weakly extracted component would then be interfered with a strong local oscillator via a second beam splitter (or coupler), enabling standard homodyne detection of a chosen quadrature.
We acknowledge the practical challenges associated with realizing this attack using current technology. Among them, the most constraining is arguably the requirement for high-bandwidth, low-noise detection—potentially in the GHz range—to resolve the short temporal modes used in state-of-the-art QKD systems. Nonetheless, the performance of photodetectors and associated readout electronics has improved significantly over the past decades, particularly in terms of bandwidth, quantum efficiency, and noise suppression. Crucially, there are no known fundamental physical limits that prevent further improvements in these areas. We thus believe the attack strategies proposed here are not only conceptually valid, but also increasingly realistic in light of technological trends.
Appendix B Analytical derivation of the QBER without attacks